Stones are thrown horizontally with the same velocity from the tops of two different buildings. One stone lands three times as far from the base of the bulilding from which it as does the other stone. Find the ratio of the height of the taller building to the height of the shorter building go pa oo 00 09 06 ㄇㄇㄇㄇㄇㄇㄇ ㄇㄇㄇㄇㄇㄧㄇ 8828823 8839
Answer #1

R1 = ux * t1
where t1 = (2h1/g)^0.5

R2 = ux * t2
where t2 = (2h2/g)^0.5

Ratio R1/R2 = t1/t2 = (h1/h2)^0.5

3/1 = (h1/h2)^0.5

h1/h2 = 9
Answer is 9